A block of mass m is observed accelerating from rest down an incline that coefficient of friction
1
Expert's answer
2021-10-03T13:22:16-0400

To find the acceleration of a block sliding down an incline, you can use the following expression:


"a=g(\sin\theta+\mu\cos\theta)."